Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probabilty Question Help >

Expert replies
by ramkarpa » Sun Aug 13, 2006 6:53 pm
Tanya prepared 4 different letters to be sent to 4 different addresses,For each letter she prepared an envelope with correct address. If the 4 letters are to be put in 4 envelopes at random, what is the probability that only 1 will be put in the correct address ?

The answer is 1/3. Any ideas any one ?

thanks
Join the discussion
Source: — Problem Solving |

solution

by viksbhat » Tue Aug 15, 2006 6:08 am
4 letters .. 4 envelopes ..
probability that first envelop gets correct letter . 1/4
probability that 2nd envelop gets wrong letter 2/3
probability that 3rd envelop gets wrong letter 1/2
probability that 3rd envelop gets wrong letter 1/1

total no of cases can happen 4 .

the probability of condition to happen is 4*1/4*2/3*1/2*1/1 = 1/3

Hope this helps.

Please feel free to comment
Join the discussion

by ajith » Thu Oct 19, 2006 12:47 am
Basically This is a Derangement problem..

Total no Ways in which u can put letters = 4! =24

according to the question we have to put 1 letter in right place

we can select this letter in 4c1 ways ie 4 ways..

We have now to put all the remaining letters in wrong envelopes this can be done 2 ways (1st letter goes to 2nd 2 to 3rd 3rd to 1st or 1st to 3rd 2nd to 1st and 3rd to 2nd , these are the only ways

so total no of ways in this condition is satisfied is 4*2=8 ways

so the probability is 8/24= 1/3

you can read more about derangements in https://en.wikipedia.org/wiki/Derangement or
https://mathworld.wolfram.com/PartialDerangement.html
Join the discussion